OpenStudy (anonymous):

find the zeros. f(x)=3x^2-6x-5 please help!

OpenStudy (anonymous):

\[\left\{x\to \frac{1}{3} \left(3-2 \sqrt{6}\right),x\to \frac{1}{3} \left(3+2 \sqrt{6}\right)\right\}\to \{x\to -0.632993,x\to 2.63299\} \]

OpenStudy (anonymous):

A plot of the function is attached.